OpenStudy (anonymous):

In a MCQ TEST comprising of 5 questions each with 4 answers. How many a student can “guess” an answer? a) 25 b) 50 c) 625 d) 120

OpenStudy (lgbasallote):

could you clarify the question?

OpenStudy (anonymous):

its 5*4! number of question* probability of choices.

OpenStudy (anonymous):

its actually an SAT question i got i thought the answer was 20 but they dont seem to think so in their choices

OpenStudy (lgbasallote):

"How many a student can 'guess' an answer" can you clarify this? some data seems to be missing

OpenStudy (anonymous):

oh yeah how many answers can a student guess sorry about that

OpenStudy (lgbasallote):

im thinking this is a combination question

OpenStudy (lgbasallote):

use \[\frac{5!}{(5-4)!}\]

OpenStudy (anonymous):

@theyatin is correct 5*4! is 120

OpenStudy (anonymous):

thank you arch its 5*4c1= 5*4!=5*4*3*2*1=120

OpenStudy (anonymous):

u mean like 5 multiplied by 4 ?

OpenStudy (lgbasallote):

\[5 \times 4! \implies 5 \times 4 \times 3 \times 2 \times 1\]

OpenStudy (anonymous):

nope 5*4! and 4!=4*3*2*1 so it will be 5*4*3*2*1=120

OpenStudy (anonymous):

oh ok yeah i got it ...

OpenStudy (lgbasallote):

that's what i wrote too anyway lol \[\frac{5!}{(5-4)!} \implies \frac{5!}{1!} \implies 5! \implies 5 \times 4 \times 3\times 2 \times 1\]

ganeshie8 (ganeshie8):

i thought its 4^5. but its not evern there in options... nvm

OpenStudy (anonymous):

is this part of algebra?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

http://en.wikipedia.org/wiki/Combination

OpenStudy (anonymous):

gee thanx p'ple

OpenStudy (lgbasallote):

wiki puts it in a complicated form but it's really just \[nCr \implies \frac{n!}{r!(n-r)!}\] lol

OpenStudy (anonymous):

right i'll try and understand it :)